# Medical-GAT: Cancer Document Classification Leveraging Graph-Based Residual Network for Scenarios with Limited Data

### v1

@@ 1 Introduction
As cancer rates rise, efficient documentation and assessment methods become vital. Electronic Health Records (EHRs) have emerged as important tools for managing vast amounts of medical data and analyzing trends, which practitioners use for optimal patient care. However, EHRs are often unstructured, with missing information and inconsistent formatting, making them challenging for cancer research. Additionally, access is restricted due to patient confidentiality, and organizations often do not disclose data for research purposes, complicating efforts to obtain the large datasets necessary to advance cancer research.
Given these limitations, there is a pressing need for alternative data sources to support cancer research. Medical abstracts, which offer insights from various publications, present a valuable resource. These abundant abstracts provide diverse perspectives on cancer types, including “thyroid”, “colon”, and “lung” cancers. To address limited data availability, we developed a dataset of 1,874 biomedical abstracts from PubMed, categorized into “thyroid cancer”, “colon cancer”, “lung cancer”, and “generic topics” (not specifically related to cancer). This dataset aims to bridge the gap left by EHR limitations and support automated cancer classification systems. Our goal is to enhance healthcare systems by providing a well-defined dataset that supports research and improves the accuracy of cancer document classification, especially where large, labeled datasets are scarce.
Building on this dataset, we implemented the R-GAT to classify cancer types using graph-based approaches. This is the first labeled and well-defined dataset that includes thyroid, colon, and lung cancers in biomedical abstracts. The R-GAT model was selected because it represents connections between entities using graphs, unlike conventional machine learning or deep learning methods that treat text as flat sequences. Its attention mechanism highlights critical information, allowing the model to gather structural context and improve performance on text analysis tasks. R-GAT is particularly beneficial in limited data scenarios, as its ability to utilize structural links enhances classification performance even with smaller datasets. Incorporating R-GAT into our dataset contributes valuable insights to cancer research and improves cancer treatment effectiveness. The key contributions of this research are outlined as follows:
@@ 2 Literature Review
Nguyen et al. [4] developed a summarization method using Dutch radiology report data, creating a hybrid model that combines an encoder-decoder language model with an attention mechanism and a Breast Imaging-Reporting and Data System (BI-RADS) score classifier. The model achieved a ROUGE-L F1-score of 51.5% for Dutch reports. Another classifier reached 83.3% accuracy, surpassing the language model’s 79.1% in BI-RADS classification. Although the model performed well, it was deemed unsuitable for clinical use, and Nguyen et al. suggested that this hybrid approach could guide future research efforts.
@@
Tang et al. [5] employed attention-based deep learning models with BERT to categorize progress notes and extract keywords. Their fine-tuned BERT model with an attention layer achieved 97.6% accuracy, outperforming the baseline. This underscores the effectiveness of hybrid and attention-based models in enhancing clinical data classification and summarization. In another study, Hepsağ et al. [6] compiled a dataset of 62 mammography records from Turkish patients, manually classified by an expert for breast cancer diagnosis. They applied conventional machine learning models, including pre-trained BERT and DistilBERT, along with an ensemble voting technique. Their results showed that BERT yielded the best performance on Turkish radiology reports, achieving a 91% F1-score.
@@
Turning to the other side, Ai et al. [7] developed the Edge-Enhanced Minimum-Margin Graph Attention Network (EMGAN) to improve short-text classification by addressing feature sparsity and weak contextual relationships, using a Heterogeneous Information Graph (HIG) and a Minimum Margin Graph Attention Network (MMGAN). Another study by Wei et al. [8] introduced a graph convolutional attention network (GCAN) to enhance the prediction of remaining useful life (RUL) in engineered systems. The method uses temporal convolution-aware nested residual connections to preserve temporal features, while an attention mechanism prioritizes the most relevant features, improving prediction accuracy.
@@
In addition, Song et al. [9] developed Graph Sequence Pre Training with Transformer (GSPT) for text-attributed graphs (TAGs), leveraging large language models (LLMs) to unify feature spaces across different graphs. GSPT emphasizes feature reconstruction to improve knowledge transferability and performance in node classification and link prediction tasks. Besides, Rao et al. [10] proposed a method for knowledge graph completion (KGC) using a Multi-layer Residual Attention Network (MRAN), which integrates categorical information with textual descriptions to improve entity embeddings. This approach enhances contextual understanding between entities and relationships, significantly improving link prediction performance.
@@
In the above studies, researchers focused on different cancers and clinical note classification using various deep learning or transformer-based approaches. Recent studies have also prioritized capturing semantic information from datasets, leading to the growing popularity of graph-based approaches in the scientific community. However, multi-cancer types such as thyroid, colon, and lung have not been extensively studied, and no publicly available datasets exist for their classification. Additionally, biomedical abstract-based text classification has received less attention, presenting an opportunity for research to develop effective solutions that can support healthcare professionals in diagnosing medical diseases or patient treatment. Furthermore, while many studies have explored graph-based techniques, the R-GAT model has not yet been utilized in previous research. However, we have addressed these gaps by creating our own dataset related to biomedical abstracts, categorizing three types of cancers. We developed a pipeline that allows effective work even with a limited data sample. Our findings and experiments will benefit the research community by providing a foundation for extending this research and developing solutions to support healthcare research.
@@ 4.2.4 Inference Testing
Figure 4: Analysis of cancer abstracts fed into the R-GAT model for classification: a) Thyroid Cancer—the model analyzed the abstract focused on the telomere-telomerase complex in both sporadic and familial thyroid cancer cases, emphasizing telomere shortening and telomerase activation; b) Lung Cancer—the R-GAT model processed an abstract detailing the effectiveness of nitrosoureas and other agents in treating various types of lung cancer, including oat cell carcinoma and adenocarcinoma. Both abstracts were correctly classified by the R-GAT model.
@@ 4.2.5 Comparative Review of Existing Studies
Additionally, examining Table 3, it is evident that transformer based models were employed in the earlier research, but the R-GAT model has not yet been applied to the classification of medical cancers. It is also discovered that previous studies were mostly concentrated on the other downstream activities indicated in Table 3, and as a result, they did not address thyroid, colon, or lung cancers.
Our objective was to implement a model capable of extracting semantic information from lengthy documents, such as medical abstracts related to thyroid, colon, and lung cancers. To validate our R-GAT model, we tested it against several transformer-based methods, including BERT, BioBERT, RoBERTa, and Bio+ClinicalBERT, as well as conventional machine learning and deep learning techniques. However, we discovered that the R-GAT is more generalizable than the other approaches we considered.
@@ 5 Conclusion and Future Work
This paper presents the first labeled dataset of biomedical abstracts related to cancers such as thyroid, colon, and lung, as well as more generic topics. The goal of this study was to assess the effectiveness of state-of-the-art methods in limited data scenarios. To achieve this, we introduced an R-GAT model for classifying cancer-related abstracts. We validated our R-GAT model through rigorous evaluation and found that it outperformed conventional machine learning (ML) and deep learning (DL) approaches in terms of generalizability. While we fine-tuned domain-specific transformer models, such as BioBERT, RoBERTa, and Bio+ClinicalBERT, their generalization performance was less effective on our curated dataset. Our findings demonstrate that the R-GAT model excels at capturing relationships between aspects in biomedical abstracts, such as diseases or treatments, by arranging the data as a graph. This strategy enables the model to take advantage of the correlations between these elements, increasing its effectiveness in circumstances with limited data. In contrast, transformer models, e.g., BioBERT, only focus on word order and may not fully capture these interactions. Moreover, transformers often require substantial datasets to function well, and when data is limited, they tend to overfit, which lowers their capacity to generalize to new or unseen data.
In future research, we aim to integrate the relational reasoning of R-GAT with the deep contextual understanding of domain-specific transformer models like BioBERT, RoBERTa, or Bio+ClinicalBERT. We envision developing a hybrid ensemble model that better captures the semantic information from the dataset and potentially improves performance. Additionally, we will look into approaches to improve the generalization capabilities of transformer-based models when used with our curated dataset.

### v3

@@ 1 Introduction
With the increasing prevalence of cancer, efficient methods for documenting and evaluating medical information have become crucial. Electronic Health Records (EHRs) have emerged as essential tools for handling large volumes of medical data and identifying patterns, which medical professionals utilize to provide optimal patient care. Nevertheless, EHRs often lack structure, contain incomplete information, and have inconsistent formatting, making them difficult to use in cancer research. Furthermore, patient privacy concerns restrict access to these records, and organizations frequently withhold data for research purposes, hindering efforts to obtain the extensive datasets necessary for advancing cancer research.
Given these constraints, there is an urgent need for alternative sources of data to support cancer research. Medical abstracts, which provide insights from various publications, offer a valuable resource. These widely available abstracts present diverse perspectives on different cancer types, including thyroid, colon, and lung cancers. To address the limited availability of data, we created a dataset comprising 1,874 biomedical abstracts from PubMed, categorized into thyroid cancer, colon cancer, lung cancer, and generic topics (not specifically related to cancer). This dataset aims to fill the gap created by EHR limitations and support automated cancer classification systems. Our objective is to enhance healthcare systems by providing a well-defined dataset that facilitates research and improves the accuracy of cancer document classification, particularly in situations where large, labeled datasets are scarce.
Utilizing this dataset, we implemented the R-GAT to classify cancer types using graph-based approaches. This represents the first labeled and well-defined dataset that includes thyroid, colon, and lung cancers in biomedical abstracts. We chose the R-GAT model because it represents connections between entities using graphs, unlike traditional machine learning or deep learning methods that treat text as linear sequences. Its attention mechanism emphasizes crucial information, enabling the model to gather structural context and enhance performance on text analysis tasks. R-GAT is particularly advantageous in scenarios with limited data, as its ability to leverage structural links improves classification performance even with smaller datasets. Incorporating R-GAT into our dataset contributes significant insights to cancer research and enhances the effectiveness of cancer treatment. The key contributions of this research are outlined as follows:
@@ 2 Literature Review
Nguyen et al. [4] created a summarization technique utilizing Dutch radiology report information. They developed a hybrid model that integrates an encoder-decoder language model with an attention mechanism and a Breast Imaging-Reporting and Data System (BI-RADS) score classifier. For Dutch reports, the model attained a ROUGE-L F1-score of 51.5%. Another classifier achieved 83.3% accuracy, outperforming the language model’s 79.1% in BI-RADS classification. Despite its strong performance, the model was infeasible for clinical application.
@@
Tang et al. [5] utilized attention-based deep learning models incorporating BERT to classify progress notes and extract key terms. Their BERT model, fine-tuned with an attention layer, reached 97.6% accuracy. This highlights the efficacy of hybrid and attention-based models in improving clinical data classification and summarization. In another study, Hepsağ et al. [6] curated a dataset comprising 62 mammography records from Turkish patients, which a subject matter expert manually classified for breast cancer diagnosis. The study also utilized state-of-the-art models, such as pre-trained BERT and DistilBERT, in addition to an ensemble voting method. However, findings from this study indicated that BERT produced the best results on Turkish radiology reports, achieving a 91% F1-score.
@@
Ai et al. [7] introduced the Edge-Enhanced Minimum-Margin Graph Attention Network (EMGAN) to address short-text classification challenges, focusing on feature sparsity and weak contextual relationships through a Heterogeneous Information Graph (HIG) and a Minimum Margin Graph Attention Network (MMGAN). Another study by Wei et al. [8] proposed a graph convolutional attention network (GCAN) for enhancing remaining useful life (RUL) predictions in engineered systems, utilizing temporal convolution-aware nested residual connections and an attention mechanism to prioritize relevant features.
@@
Song et al. [9] developed Graph Sequence Pre Training with Transformer (GSPT) for text-attributed graphs (TAGs), leveraging large language models (LLMs) to unify feature spaces across diverse graphs, emphasizing feature reconstruction to boost knowledge transferability and performance in node classification and link prediction tasks. Meanwhile, Rao et al. [10] presented a Multi-layer Residual Attention Network (MRAN) for knowledge graph completion (KGC), integrating categorical information with textual descriptions to enhance entity embeddings and improve contextual understanding between entities and relationships.
@@
These studies explored various cancers and clinical note classification using deep learning and transformer-based approaches, with a growing emphasis on capturing semantic information through graph-based methods. However, multi-cancer types such as thyroid, colon, and lung remain understudied, with no publicly available datasets for their classification. Biomedical abstract-based text classification has also received limited attention, presenting an opportunity for developing solutions to support healthcare professionals in disease diagnosis and treatment planning. While graph-based techniques have been widely explored, the R-GAT model has not been previously utilized in this context. To address these gaps, we created a dataset of biomedical abstracts categorizing three cancer types and developed a pipeline effective for limited data samples. Our research contributes to the field by providing a foundation for future investigations and supporting healthcare research through innovative solutions. Our findings will help the research community discover creative solutions to healthcare issues and progress medical research.
@@ 4.2.4 Inference Testing
Figure 4: Analysis of cancer abstracts fed into the R-GAT model for classification: a) Thyroid Cancer—the model analyzed the abstract focused on the telomere-telomerase complex in both sporadic and familial thyroid cancer cases, emphasizing telomere shortening and telomerase activation; b) Lung Cancer—the R-GAT model processed an abstract detailing the effectiveness of nitrosoureas and other agents in treating various types of lung cancer, including oat cell carcinoma and magenta adenocarcinoma. Both abstracts were correctly classified by the R-GAT model.
@@ 4.2.5 Comparative Review of Existing Studies
Additionally, examining Table 3, it is evident that transformer-based models were employed in the earlier research, but the R-GAT model has not yet been applied to classify medical cancers. It is also discovered that previous studies were mostly concentrated on the other downstream activities indicated in Table 3, and as a result, they did not address thyroid, colon, or lung cancers.
We aimed to implement a model capable of extracting semantic information from lengthy documents, such as medical abstracts related to thyroid, colon, and lung cancers. To validate our R-GAT model, we tested it against several transformer-based methods, including BERT, BioBERT, RoBERTa, and Bio+ClinicalBERT, as well as conventional machine learning and deep learning techniques. However, we discovered that the R-GAT is more generalizable than the other approaches we considered.
@@ 5 Conclusion and Future Work
This study introduces a labeled dataset of biomedical abstracts covering specific cancer types (thyroid, colon, and lung) and broader topics. The research aimed to evaluate the performance of cutting-edge techniques in scenarios with limited data availability. To this end, we developed an R-GAT model for categorizing cancer-related abstracts. Our R-GAT model underwent thorough testing and was found to surpass traditional machine learning (ML) and deep learning (DL) methods in terms of generalization capabilities. Although we optimized domain-specific transformer models, including BioBERT, RoBERTa, and Bio+ClinicalBERT, their ability to generalize effectively on our curated dataset was less impressive. Our results indicate that the R-GAT model is particularly adept at identifying relationships between various aspects of biomedical abstracts, such as diseases or treatments, by structuring the data as a graph. This approach allows the model to leverage correlations between these elements, enhancing its efficacy in limited data. In contrast, transformer models like BioBERT primarily consider word sequences and may not fully capture these interconnections. Furthermore, transformers often require large datasets to work effectively, and in data-scarce contexts, they are prone to overfitting, limiting their ability to generalize to new or unknown information. To this end, this study adopt a solution that can mitigate the issue of the domain specific transformer such as BIO+BERT or RoBertA.
Our future research plans involve combining the relational reasoning capabilities of R-GAT with the advanced contextual understanding provided by domain-specific transformer models such as BioBERT, RoBERTa, or Bio+ClinicalBERT. We anticipate creating a hybrid ensemble model that enhances the extraction of semantic information from the dataset, potentially leading to improved performance. Furthermore, we will explore methods to enhance the generalizability of transformer-based models when applied to our specially curated dataset.